• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:01:01 on April27,2004

    Entry number 119465

    keyword=Swing Shift Summary

    Swing shift, 4/26/2004
    Shift workers: H. Benaoum, J. Reinhold.

    Program is to continue production run with Be target.
    16:00 Shift started with no beam, Problem with power supply in SL.
    16:00- 16:30 CODA problem from previous shift: begin-run failed. ROC16 did not come back, start-up values were gone. Called Bob, Bob did something fixed it. Restarted coda afew times. Problem with ROC1,
    after reboot, all is fine now. Waiting for beam.
    16:35 Called MCC and was told that they fixed the RF power supply problem and are getting ready to get beam back to us (in 10 minutes?)
    16:54 MCC called to inform us that PD decided to run a test plan before sending beam back to us. It will be 15 minutes.
    17:00 beam back, checked raster size.
    Run 2803 started at 17:03, Beam was not stable for the first ten minutes: there was a large shift in energy split measurements.
    Run ended at 17:39, 1M events.
    Run 2804 started at 17:40, ended at 18:18, 1M
    Run 2805 started at 18:18:19, ended at 18:51, 1M
    Bodo took over to get ready for a e-p energy measurement.
    MCC tune beam for Hall C. changed back to data taking.
    Run 2806 started at 19:09, Rich tripped. Reset. stopped at 19:15, 94K.
    Run 2807 started at 19:15, Stopped at 19:45, 1M events.
    19:45 Bodo tookover to do e-p. Moved to Empty.
    21:00 Body completed e-p, we moved tragte back to Be.
    Asked for fast raster back, originial beam position and 1C12 OTR in.
    21:15 started run 2808, edned at 21:45, 1M
    21:45 started run 2809, ended at 22:18, 1M
    22:18 started run 2810, ended at 22:51, 1M
    22:55 started run 2811, ended at, RICH HV tripped,
    stopped run at 23:04, with 313K, Reset RICH HV.
    23:10 started run 2812, ended at 23:46, 1M
    23:47 started run 2813, still ongoing at the end of the shift.
    24:00 change to next shift with good beam 100 uA.